# Perfect Quantum State Revivals: Designing Arbitrary Potentials with Specified Energy Levels

### v1

@@ IV.3 Potentials with alternating gaps between levels
As the next example, let us construct a potential with the gap between the subsequent levels alternating betwen two values. Starting from the harmonic oscillator, we construct $N=100$ additional levels following this pattern.
@@ IV.4 Potentials with prime and Fibonacci spectra
V Multidimensional case
@@
VI Conclusions
We have presented a general method for desiging potentials that exhibit perfect quantum state revivals. The examples given demonstrate the power and flexibility of the method; there exist endless possibilities for other energy level patterns that obey Rule (2); constructing their corresponding quantum potentials leads to a multitude of potentials that exhibit perfect wavepacket revivals. The intertwining method described above and applied in Algorithm 1 can also be used to quickly and easily design potentials that have any desired energy level positions and is thus a powerful inverse design technique that may possibly have applications beyond what we have considered here. Since separation of variables allows the construction of $n$-dimensional potentials from individual 1-dimensional potentials, the spectral engineering of any desired potential is conceivable with the algorithm presented. A possible extension of our work could include designing potentials for relativistic quantum revivals [5].

### v2

@@ IV.3 Potentials with alternating gaps between levels
As the next example, let us construct a potential with the gap between the subsequent levels alternating between two values. Starting from the harmonic oscillator, we construct $N=100$ additional levels following this pattern.
@@ IV.4 Potentials with prime and Fibonacci spectra
V Error analysis
Naturally, it is of interest to know the level of precision with which the actual energy levels of the potentials obtained by our method match the desired levels used to design them. We have performed such analysis for the six potentials discussed in the previous section. To calculate the energy spectrum, we have implemented the finite element method where the maximum cell size (the variable “MaxCellMeasure” in the software Mathematica) can be chosen. For various values of this variable we have then calculated the energy spectrum, compared it with the desired energy levels, and evaluated the maximum error. The results are shown in Fig. 8. The best results were obtained for the reverse biperiodic and biperiodic potentials while the worse ones were obtained for the prime and Fibonacci spectra. As a general rule, in the log-log scale all the curves have a slope of approximately 2, meaning that decreasing the maximum cell size by an order of magnitude results in the maximum error reduction by two orders of magnitude. This behavior suggests that the errors are not caused so much by the potentials themselves, but rather by the procedure of calculating the levels. This way, our method enables to design potentials that have the desired energy levels with a high degree of precision.
Figure 8: The dependence of maximum errors of actual energy levels on the maximum cell size used in the finite element method for the particular potentials discussed in Sec. IV.
VI Multidimensional case
@@
VII Conclusions
We have presented a general method for designing potentials that exhibit perfect quantum state revivals. The examples given demonstrate the power and flexibility of the method; there exist endless possibilities for other energy level patterns that obey Rule (2); constructing their corresponding quantum potentials leads to a multitude of potentials that exhibit perfect wavepacket revivals. The intertwining method described above and applied in Algorithm 1 can also be used to quickly and easily design potentials that have any desired energy level positions and is thus a powerful inverse design technique that may possibly have applications beyond what we have considered here. Since separation of variables allows the construction of $n$-dimensional potentials from individual 1-dimensional potentials, the spectral engineering of any desired potential is conceivable with the algorithm presented. Finally, it should be pointed out that the operators $A^{+}$ and $A^{-}$ are not unique [20]; higher-order operators would allow isospectral deformations similar to what we have shown, but with additional freedom [26, 27, 28]. Along with consideration of higher-order intertwining operators, a possible extension of our work could include designing potentials for relativistic quantum revivals [5].